Let u = 31 - ), v= 41+j, w = i +5j Find the specified scalar. (4u) .v (4u)•v=0 Enter your answer in the answer box
Answer #1

Given u=3i-j v- yit; (4u.v) = 4[u.v] = 4[ (x4) – (1+1)] 4[12-1] = 44. (4.v) = (121-45) • (uits) us- 4 = 44. -